OpenStudy (anonymous):

Which shows the equation below written in the form ax2 + bx + c = 0?

OpenStudy (anonymous):

x2 + 6x + 8 = 3

OpenStudy (anonymous):

do I just subtract 3 from 8?

OpenStudy (anonymous):

so would it be x^2 + 6x + 5 = 0?

OpenStudy (aum):

correct.

OpenStudy (anonymous):

so for this one what would be my first step to converting it? 2x^2-3=-4x-1

OpenStudy (anonymous):

2x^2-3=-4x-1

OpenStudy (aum):

We have to make the right hand side of the equation zero. Therefore, Add 4x to both sides. Then add 1 to both sides.

OpenStudy (anonymous):

Alright so it should be 2x^2 + 4x - 2 = 0 Thanks for the help